OpenStudy (brelove13):

a right triangle has an area of 20 square inches. The square of the hypotenuse is 116. find the lengths pf the legs pf the triangle.

OpenStudy (marigirl):

is it correct to use the base measurement as sqrt166 which is 12.88

OpenStudy (brelove13):

it's actually 116

OpenStudy (agent0smith):

It's 116 isn't it? The hypotenuse squared. So \[\large 116 = base^2+height ^2\]

OpenStudy (marigirl):

sorry typo.

OpenStudy (brelove13):

@Kkutie7 what happened to your equation?

OpenStudy (kkutie7):

humm?

OpenStudy (kkutie7):

oh I deleted it! I'm sorry. I'm kind of out of it right now. I basically solved for one component and substituted into the other equation.

OpenStudy (brelove13):

\[a ^{2}+b ^{2}=116\]

OpenStudy (kkutie7):

\[116=base^{2}+height^{2}\] \[20=\frac{1}{2}base*height\]

OpenStudy (brelove13):

\[(\frac{ 40 }{ b})^{2}+b ^{2}=116\]

OpenStudy (brelove13):

20=\[\frac{ 1 }{2 }base \times height\] 40=\[base \times height\]

OpenStudy (brelove13):

\[\frac{ 40 }{ base }=height\]

OpenStudy (brelove13):

and then i sub the height from this equation into the pythagorean theorem equation?

OpenStudy (brelove13):

\[116=\left(\begin{matrix}40 \\ base\end{matrix}\right)^{2}+b ^{2}\]

OpenStudy (brelove13):

but what do i do next?

OpenStudy (agent0smith):

Multiply both sides (every term) by b^2